OpenStudy (anonymous):

What is the slope of this line? y = -4x + 7 A. 4 B. 7 C. -4 D. -7

pooja195 (pooja195):

y=mx+b m=slope

OpenStudy (anonymous):

4

OpenStudy (anonymous):

lel jk

OpenStudy (texaschic101):

it is not 4

OpenStudy (anonymous):

its -4

OpenStudy (texaschic101):

yep ^^

OpenStudy (anonymous):

Thx u guys

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

y=mx=b m is your slope and b is your y-intercept

OpenStudy (anonymous):

can u help with me with more?

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

What is the y-intercept of this line? y = 5x + 8

OpenStudy (anonymous):

8 because the back number without the x will always be your y intercept. the numbers with the x next to them are your slope

OpenStudy (anonymous):

What is the equation of the given line in Standard Form? A. 2x + y = 4 B. y = 2x + 4 C. y - 4 = 2x D. -2x + y = 4

OpenStudy (anonymous):

What is the equation of the given line in Standard Form? A. 2x + y = 4 B. y = 2x + 4 C. y - 4 = 2x D. -2x + y = 4

OpenStudy (anonymous):

well which do you think it is? standard form is y=Mx+b which choice is set up that way?